The rms speed of the gas at 300 K is 1370 m/s. What is its rms speed at 1200 K? Numeric Answer:
Answer #1

rms speed at 1200 K = 2740 m/s

Explanation

rms speed of a gas is directly proportional to square root of temperature

rms2 / rms1 = (T2 / T1)1/2

where

rms1 = rms speed at 300 K = 1370 m/s

T1 = 300 K

rms2 = rms speed at 1200 K

T2 = 1200 K

Substituting the values

rms2 / (1370 m/s) = (1200 K / 300 K)1/2

rms2 / (1370 m/s) = (4)1/2

rms2 / (1370 m/s) = 2

rms2 = (1370 m/s) * 2

rms2 = 2740 m/s